OpenStudy (anonymous):

6. Which ordered pair is a solution of the equation y = 3x? (1 point) (–2, –9) (–8, –18) (–8, –3) (–10, –30)

OpenStudy (anonymous):

10. w – 2 = –3 (1 point) w = –3/2 w = –5 w = –1 w = 6

OpenStudy (anonymous):

11. 2.8 = 2y (1 point)y = 1.4 y =1/2 y = 2.8 y = 0.6